# Thermodynamic and energetic constraints on out-of-equilibrium tunneling rates

### v1

@@ Appendix D Derivation of the thermodynamic constraint
where we recall that $\hat{A}_{\text{H}}(0)=\hat{A}$.
Appendix E Grönwall inequality on out-of-equilibrium rates
@@
From the thermodynamic constraint in Eq. (16), it is possible to remove the out-of-equilibrium response of tunneling rates, $\frac{\partial\Gamma_{\rightleftarrows}}{\partial\beta_{\text{L}}}(\beta_{\text{L}},\beta_{\text{R}})$, by means of Grönwall’s lemma [76]. Indeed, Eq. (16) can be cast as
@@
$\begin{split}y(x)&\equiv\Gamma_{\rightleftarrows}(x,\beta_{\text{R}}),\\ f(x)&\equiv\frac{\partial\Gamma_{\rightleftarrows}}{\partial\beta_{\text{L}}}(\beta_{\text{R}},\beta_{\text{R}})-\Delta F_{\text{L}}^{\text{(c)}}(x)\eta^{\text{(h)}}(x)\Gamma_{\rightleftarrows}(\beta_{\text{R}},\beta_{\text{R}}),\\ g(x)&\equiv-\Delta F_{\text{L}}^{\text{(h)}}(x)\eta^{\text{(c)}}(x).\end{split}$ (71)
@@
we see that $v(\beta_{\text{R}})=1$ and $v(x)\geq 0$ for all $x$. The derivative of $y(x)/v(x)$ reads
@@
Integrating this expression from $\beta_{\text{L}}$ to $\beta_{\text{R}}$ leads to
@@
Appendix F Derivation of the energetic constraint
In this Appendix, we prove the energetic constraint in Eq. (20) in differential form, namely
$\frac{\partial^{2}\Gamma_{\rightleftarrows}}{\partial\beta_{\text{L}}^{2}}(\beta_{\text{L}},\beta_{\text{R}})\geq\frac{\partial U_{\text{L}}}{\partial\beta_{\text{L}}}(\beta_{\text{L}})\Gamma_{\rightleftarrows}(\beta_{\text{L}},\beta_{\text{R}}).$ (76)
By integrating over $\beta_{\text{L}}$ from an initial inverse temperature $\beta_{\text{L,1}}$ to a final one $\beta_{\text{L,2}}$, one can then reach the integral form in Eq. (20). As in Appendix D, here we prove the energetic constraint only for $\Gamma_{\rightarrow}$ since the proof for $\Gamma_{\leftarrow}$ is analogous. First, notice that the $\beta_{\text{L}}$-derivatives act only on the occupation probabilities of subsystem L,
$\frac{\partial^{2}\Gamma_{\rightarrow}}{\partial\beta_{\text{L}}^{2}}=\frac{4\pi}{\hbar}\sum_{nmlk}|\mathcal{A}_{nmlk}|^{2}\delta(\epsilon_{mn}^{\text{(L)}}+\epsilon_{kl}^{\text{(R)}}-\hbar\omega)\frac{\partial^{2}p_{n}^{\text{(L)}}}{\partial\beta_{\text{L}}^{2}}p_{l}^{\text{(R)}}.$ (77)
Then, recalling that the internal energy satisfies $U_{\text{L}}(\beta_{\text{L}})=-\frac{\partial\log Z_{\text{L}}(\beta_{\text{L}})}{\partial\beta_{\text{L}}}$, the derivative of the occupation probabilities reads
$\begin{split}\frac{\partial^{2}p_{n}^{\text{(L)}}}{\partial\beta_{\text{L}}^{2}}&=\frac{\partial}{\partial\beta_{\text{L}}}\left[\left(-\epsilon_{n}^{\text{(L)}}+U_{\text{L}}(\beta_{\text{L}})\right)p_{n}^{\text{(L)}}\right]\\ &=\left[\frac{\partial U_{\text{L}}(\beta_{\text{L}})}{\partial\beta_{\text{L}}}+\left(U_{\text{L}}(\beta_{\text{L}})-\epsilon_{n}^{\text{(L)}}\right)^{2}\right]p_{n}^{\text{(L)}}\\ &\geq\frac{\partial U_{\text{L}}(\beta_{\text{L}})}{\partial\beta_{\text{L}}}p_{n}^{\text{(L)}}.\end{split}$ (78)
Substituting this inequality in Eq. (77) we find Eq. (76). Furthermore, we see here that the inequality is saturated when $|\mathcal{A}_{nmlk}|^{2}\delta(\epsilon_{mn}^{\text{(L)}}+\epsilon_{kl}^{\text{(R)}}-\hbar\omega)$ is nonzero only for energies $\epsilon_{n}^{\text{(L)}}$ close to the internal energy $U_{\text{L}}(\beta_{\text{L}})$.
Appendix G Sufficient condition for trivial constraints
Both thermodynamic and energetic cost, Eqs. (15) and (20) respectively, are positive by definition, and are lower-bounded by the rate response $\mathcal{W}^{\text{(Resp)}}$ in Eq. (17). Therefore, the thermodynamic and energetic constraints are nontrivial whenever $\mathcal{W}_{\rightleftarrows}^{\text{(Resp)}}>0$. In this appendix we provide sufficient conditions for $\mathcal{W}^{\text{(Resp)}}_{\rightleftarrows}\lessgtr 0$, making the results of Eqs. (16, 20) trivial. Like the previous appendices, we provide the derivation for the forward process. Starting from the definition of the rate response, the derivatives with respect to $\beta_{\text{L}}$ act only on the probabilities $p_{n}^{\text{(L)}}$. Specifically, the rate response becomes
$\begin{split}\mathcal{W}^{\text{(Resp)}}_{\rightarrow}&=\frac{2\pi}{\hbar}\sum_{nmlk}|\mathcal{A}_{nmlk}|^{2}\delta(\epsilon_{mn}^{\text{(L)}}+\epsilon_{kl}^{\text{(R)}}-\hbar\omega)\times\\ &\times\left[\partial_{\text{L}}p_{n}^{\text{(L)}}(\beta_{\text{L,2}})-\partial_{\text{L}}p_{n}^{\text{(L)}}(\beta_{\text{L,1}})\right]p_{l}^{\text{(R)}}(\beta_{\text{R}})\end{split}$ (79)
where the derivative of the probability $p_{n}^{\text{(L)}}$ reads
$\partial_{\text{L}}p_{n}^{\text{(L)}}(\beta)=\left[U_{\text{L}}(\beta)-\epsilon_{n}^{\text{(L)}}\right]p_{n}^{\text{(L)}}(\beta).$ (80)
Taking a second derivative with respect to $\beta$ we find
$\partial_{\text{L}}^{2}p_{n}^{\text{(L)}}(\beta)=\left[\partial_{\text{L}}U_{\text{L}}(\beta)+[U_{\text{L}}(\beta)-\epsilon_{n}^{\text{(L)}}]^{2}\right]p_{n}^{\text{(L)}}(\beta).$ (81)
We now focus on all energies $\epsilon_{n}^{\text{(L)}}$ lying far away from the internal energy $U_{\text{L}}$. Namely, we suppose that the following inequality holds for all $\beta\in[\beta_{\mathrm{L},2},\beta_{\mathrm{L},1}]$,
$|U_{\text{L}}(\beta)-\epsilon_{n}^{\text{(L)}}|\geq\sqrt{-\partial_{\text{L}}U_{\text{L}}(\beta)}=\delta U_{\text{L}}(\beta),$ (82)
where $\delta U_{\text{L}}(\beta)=\sqrt{\sum_{n}(\epsilon_{n}^{\text{(L)}})^{2}p_{n}^{\text{(L)}}(\beta)-U_{\text{L}}^{2}(\beta)}$ is the standard deviation of the internal energy. In this energy interval, the second derivative of the occupation probabilities is positive, i.e. $\partial^{2}_{\text{L}}p_{n}^{\text{(L)}}(\beta)\geq 0$, making $\partial_{\text{L}}p_{n}^{\text{(L)}}(\beta)$ an increasing function in the inverse temperature $\beta$. This in turn makes the rate response negative,
$\partial_{\text{L}}p_{n}^{\text{(L)}}(\beta_{\text{L,1}})-\partial_{\text{L}}p_{n}^{\text{(L)}}(\beta_{\text{L,2}})\geq 0\quad\Rightarrow\quad\mathcal{W}^{\text{(Resp)}}_{\rightarrow}\leq 0,$ (83)
thus trivializing both thermodynamic and energetic constraint.
Conversely, the initial states’ energies lying within a standard deviation from the internal energy for all inverse temperatures in $[\beta_{\text{L,2}},\beta_{\text{L,1}}]$ is a sufficient condition for the rate response to be positive $\mathcal{W}^{\text{(Resp)}}_{\rightleftarrows}\geq 0$. Indeed, this condition on the energies makes $\partial_{\text{L}}p_{n}^{\text{(L)}}(\beta)$ a decreasing function in the inverse temperature $\beta$, leading to
$\partial_{\text{L}}p_{n}^{\text{(L)}}(\beta_{\text{L,1}})-\partial_{\text{L}}p_{n}^{\text{(L)}}(\beta_{\text{L,2}})\leq 0\quad\Rightarrow\quad\mathcal{W}^{\text{(Resp)}}_{\rightarrow}\geq 0.$ (84)
Specifically, for the two level system considered in Sec. IV.1, we can easily compute both the internal energy and its standard deviation

### v2

@@ Appendix C Transition rates in the weak-tunneling regime
$\hat{\mathcal{U}}_{0}(t,0)$ $\equiv e^{-i\hat{H}_{0}t/\hbar},$ (58a)
[…]
are the evolution induced by the free Hamiltonian $\hat{H}_{0}$ and the first correction due to the tunneling Hamiltonian $\hat{V}(t)$, see Eq. (20). For the conditional probability between two common eigenstates $a,b$ of the Hamiltonian $\hat{H}_{0}$, this means
$|\braket{b}{\mathcal{U}(t,0)}{a}|^{2}\approx|\langle b|\hat{\mathcal{U}}_{0}(t,0)+\delta\hat{\mathcal{U}}(t,0)|a\rangle|^{2}$ (59)
$\approx\left|\delta_{ab}-i\left[A_{ba}\frac{e^{-i\omega t}-e^{-i\epsilon_{ba}t/\hbar}}{\epsilon_{ba}-\hbar\omega}+A_{ab}^{*}\frac{e^{i\omega t}-e^{-i\epsilon_{ba}t/\hbar}}{\epsilon_{ba}+\hbar\omega}\right]\right|^{2}$
where we used the matrix elements of the tunneling Hamiltonian $A_{ba}=\langle b|\hat{A}|a\rangle$ and the energy differences $\epsilon_{ba}=\epsilon_{b}-\epsilon_{a}$. If we now further assume that the tunneling operator allows for either the transition $|a\rangle\to|b\rangle$ or $|b\rangle\to|a\rangle$ but not for both, i.e. $A_{ab}A_{ba}=0$, we find
$\begin{split}|\braket{b}{\mathcal{U}(t,0)}{a}|^{2}&\approx|A_{ba}|^{2}\frac{2-2\cos\left[\left(\epsilon_{ba}/\hbar-\omega\right)t\right]}{\left(\epsilon_{ba}-\hbar\omega\right)^{2}}\\ &+|A_{ab}|^{2}\frac{2-2\cos\left[\left(\epsilon_{ba}/\hbar+\omega\right)t\right]}{\left(\epsilon_{ba}+\hbar\omega\right)^{2}}\,,\end{split}$ (60)
for $b\neq a$. The long-time limit tunneling rates from state $a$ to state $b$ are then simply obtained by taking a time derivative, together with the limit $t\to\infty$, while keeping $\frac{|A_{ba}|t}{\hbar},\frac{|A_{ab}|t}{\hbar}\ll 1$,
$\partial_{t}|\braket{b}{\mathcal{U}(t,0)}{a}|^{2}\xrightarrow{t\to\infty}$
$\frac{2\pi}{\hbar}|A_{ba}|^{2}\delta(\epsilon_{ba}-\hbar\omega)+\frac{2\pi}{\hbar}|A_{ab}|^{2}\delta(\epsilon_{ba}+\hbar\omega)\ .$ (61)
We now choose the transition matrix element $A_{ba}$ to be the non-zero one. The transition rates are obtained by multiplying by the probability of the initial state
$\Gamma_{a\to b}$ $=$ $\frac{2\pi}{\hbar}\delta(\epsilon_{ba}-\hbar\omega)|A_{ba}|^{2}p_{a},$ (62)
$\Gamma_{a\leftarrow b}$ $=$ $\frac{2\pi}{\hbar}\delta(\epsilon_{ba}-\hbar\omega)|A_{ba}|^{2}p_{b}.$ (63)
Finally, to get the full transport rates, we multiply by the probability of the initial state and sum over all initial and final states contributing to the process of absorption or emission of $\hbar\omega$, namely $\Gamma_{\rightarrow}=\sum_{a,b}\Gamma_{a\to b}$ and $\Gamma_{\leftarrow}=\sum_{a,b}\Gamma_{a\leftarrow b}$. Specializing to a bipartite system, with $\hat{H}_{0}=\hat{H}_{\text{L}}+\hat{H}_{\text{R}}$ and $\hat{H}_{\text{L}}\ket{nl}=\epsilon_{n}^{(\text{L})}\ket{nl}$, $\hat{H}_{\text{R}}\ket{nl}=\epsilon_{l}^{(\text{R})}\ket{nl}$, the rates take the form
$\begin{split}\Gamma_{mk\leftarrow nl}&\equiv\frac{2\pi}{\hbar}\delta(\epsilon_{mn}^{\text{(L)}}+\epsilon_{kl}^{\text{(R)}}-\hbar\omega)|A_{mk,nl}|^{2}p_{l}^{\text{(R)}}p_{n}^{\text{(L)}},\\ \Gamma_{mk\to nl}&\equiv\frac{2\pi}{\hbar}\delta(\epsilon_{mn}^{\text{(L)}}+\epsilon_{kl}^{\text{(R)}}-\hbar\omega)|A_{mk,nl}|^{2}p_{k}^{\text{(R)}}p_{m}^{\text{(L)}},\end{split}$ (64)
leading to Eq. (21) in the main text.
Appendix D Current and noise in the weak-tunneling regime
It is interesting to connect these rates to transport quantities. On one hand, this allows us to exploit our constraints to develop bounds on the noise under a temperature bias in terms of currents. On the other hand, it provides an experimental access to test our bounds on transition probabilities and rates.
We start by defining a current operator via Heisenberg’s equation of motion as
$\hat{I}(t)\equiv-\frac{i}{\hbar}\left([\hat{Q},\hat{A}]e^{-i\omega t}+[\hat{Q},\hat{A}^{\dagger}]e^{i\omega t}\right),$ (65)
[…]
With Eq. (58), the average current can be expanded in powers of the tunneling operator $A$ as $\langle\hat{I}_{\text{H}}(t)\rangle\approx I^{(1)}(t)+I^{(2)}(t)$ with
$I^{(2)}(t)$ $\equiv\text{Tr}\left\{[\delta\mathcal{U}^{\dagger}(t,0)\hat{I}(t)\hat{\mathcal{U}}_{0}(t,0)+\hat{\mathcal{U}}_{0}^{\dagger}(t,0)\hat{I}(t)\delta\mathcal{U}(t,0)]\hat{\rho}_{0}\right\},$ (66b)
with $\hat{I}(t)$ given in Eq. (65) and $\hat{\rho}_{0}$ is the state of the system at time $t=0$. Connecting to Appendix C, we assume that $[\hat{\rho}_{0},\hat{H}_{0}]=0$, such that, if {|$\{\ket{a}\}$ are the eigenstates of $\hat{H}_{0}$ with energies $\epsilon_{a}$, their occupations are given by $\hat{\rho}_{0}\ket{a}=p_{a}\ket{a}$. For the zero-frequency noise, instead we start from the expansion
$\begin{split}\mathcal{S}(t)&\approx\langle\hat{I}_{\text{H}}(t)\hat{I}(0)\rangle\approx\text{Tr}\left\{\hat{\mathcal{U}}_{0}^{\dagger}(t,0)\hat{I}(t)\hat{\mathcal{U}}_{0}(t,0)\hat{I}(0)\hat{\rho}_{0}\right\}.\end{split}$ (67)
We now again assume that the tunneling operator satisfies $A_{ab}A_{ba}=0$, which means that only one of the transitions between |$\ket{a}\to\ket{b}$ and |$\ket{b}\to\ket{a}$ is possible. After some algebra, we find the current and noise expressed in terms of the transition rates (62) and (63)
$\begin{split}I&=\sum_{ab}(q_{b}-q_{a})[\Gamma_{b\to a}-\Gamma_{b\leftarrow a}],\\ \mathcal{S}&=\sum_{ab}(q_{b}-q_{a})^{2}[\Gamma_{b\to a}+\Gamma_{b\leftarrow a}],\\ \end{split}$ (68)
where we recall that $q_{a}\ket{a}=\hat{Q}\ket{a}$. If we now add the hypothesis $[\hat{Q},\hat{A}]=q\hat{A}$ as for standard charge currents, such that one has $(q_{b}-q_{a})\to q$, the results for current and noise in terms of rates of Sec. III in the main text are thus recovered.
Appendix E Grönwall inequality on out-of-equilibrium rates in a bipartite system
@@
From the thermodynamic constraint in Eq. (23a), it is possible to remove the out-of-equilibrium response of tunneling rates, $\partial_{\text{L}}\Gamma_{\leftrightarrows}(\beta_{\mathrm{h}},\beta_{\mathrm{c}})$, by means of Grönwall’s lemma [63]. Indeed, Eq. (23a) can be cast as
@@
$\begin{split}y(x)&\equiv\Gamma_{\rightleftarrows}(x,\beta_{\mathrm{c}}),\\ f(x)&\equiv\partial_{\text{L}}\Gamma_{\leftrightarrows}(\beta_{\mathrm{c}},\beta_{\mathrm{c}})-\Delta F_{\text{L}}^{\text{(c)}}(x)\eta^{\text{(h)}}(x)\Gamma_{\rightleftarrows}(\beta_{\mathrm{c}},\beta_{\mathrm{c}}),\\ g(x)&\equiv-\Delta F_{\text{L}}^{\text{(h)}}(x)\eta^{\text{(c)}}(x).\end{split}$ (70)
@@
we see that $v(\beta_{\mathrm{c}})=1$ and $v(x)\geq 0$ for all $x$. The derivative of $y(x)/v(x)$ reads
@@
Integrating this expression from $\beta_{\mathrm{h}}$ to $\beta_{\mathrm{c}}$ leads to